OpenStudy (mayaal):

i will give u a medal and a fan!!!!!!!!! if tan(theta)=-1,and the sin(theta)>0,what is the value of cos(theta)?

OpenStudy (mayaal):

@ganeshie8 ????????

OpenStudy (mathmate):

Use the identity \(\large {tan(\theta) = \frac{sin(\theta)}{cos(\theta)}}\)

OpenStudy (mayaal):

ok,then what?

OpenStudy (mayaal):

@mathmate ?

OpenStudy (mathmate):

Oh, I thought they gave you the value of sin(theta). Sorry, now use this: \[sec^2(\theta)=1+tan^2(\theta)\] but since \(sec(\theta)=\frac{1}{cos(\theta)}\) you will be able to solve for cos.

OpenStudy (mayaal):

and then what do i do after solving for cos?

OpenStudy (mathmate):

That seems to be the answer the question is asking for... unless I'm mistaken again! :)

OpenStudy (mayaal):

let me give u the solution of the question:

OpenStudy (mayaal):

Since tan(theta) < 0 and sin(theta) > 0, the terminating side of the angle is in the second quadrant. tan (theta)= -y/x = 1/-1 = -11; r = square root of (1 ^2 + 1 ^2) = square root of 2; cos (theta) = negative one over the square root of 2 = negative (square root of 2 over 2)

OpenStudy (mayaal):

i know this is really confusing!

OpenStudy (mathmate):

Not at all!

OpenStudy (mayaal):

how did v get to this solution?can u plz help me?

OpenStudy (mathmate):

First, you've followed the right steps. tan(theta)=-y/x=-1

OpenStudy (mathmate):

Since tan(theta)=sin(theta)/cos(theta), AND sin(theta)>0

OpenStudy (mathmate):

cos(theta) must be <0. ok so far?

OpenStudy (mayaal):

yes:)

OpenStudy (mathmate):

Also, since tan(theta)=-1, so cos(theta)=-sin(theta), ok there?

OpenStudy (mayaal):

yes,u solved for cos theta,right?

OpenStudy (mathmate):

Yes, all you need is after solving for cos(theta), put a negative sign on it, since sin(theta)>0, and cos(theta)=-sin(theta), so cos(theta)<0.

OpenStudy (mayaal):

yep,but then how did they find the radius,that is,square root of 2?

OpenStudy (mathmate):

As you did earlier, you had sec^2(theta)=2 so \(cos(\theta)=\pm \sqrt{2}/2\) and we know that it is negative.

OpenStudy (mathmate):

Use the identity \[sec^2(\theta)=1+tan^2(\theta)=1+(-1)^2=2\] so \[cos^2(\theta)=\frac{1}{2}\]

OpenStudy (mayaal):

oh,ok.so to solve for cos,we will find the square root of 1/2,right

OpenStudy (mathmate):

Yep.

OpenStudy (mayaal):

so that's basically it?

OpenStudy (mathmate):

Note also that after rationalizing the denominator: \[\sqrt{\frac{1}{2}}=\frac{\sqrt2}{2}\]

OpenStudy (mayaal):

yes,i get it now.Thankyou sooooooooooo much @mathmate !

OpenStudy (mathmate):

Yes, that's all to it. When problems are broken down to pieces, they always seem easy! :)

OpenStudy (mayaal):

that's right:)

OpenStudy (mathmate):

You're welcome! @mayaal !